A financial planner recommends a plan for a client who has $64,000 to invest. (See the chart.) At the end of the presentation, the client asks, "How much will be invested at each rate?" Answer this question using P = 64,000, r1 = 14, r2 = 6.4, and I = 7,310.80.
$___________ at 14%
$___________ at 6.4%

Answer by Boreal(15235)

x=amount at 14%
64000-x= amount at 6.4%
.14x+(.064)(64000-x)=7310.80
.14x-.064x+4096=7310.80
0.076x=3214.80
x=$42300@14%=$5922
64000-x=$21700@0.064=$1388.80
Add to $7310.80